# On Mirzakhani’s twist torus conjecture

### v2

@@ 5.3. Orbit closures of twist tori
Let $\mathcal{L}=\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})}$. If $\varepsilon=-1$ and
$g+\frac{1}{4}\#\underline{\kappa}_{\operatorname{odd}}+\frac{1}{2}\leq\#\gamma-N_{co},$
then $\mathcal{L}$ is the ambient component of $\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};\varepsilon)$ containing $\mathcal{L}$.
@@
That $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})\subset\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};\varepsilon)$, hence that $\mathcal{L}\subset\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};\varepsilon)$, follows from the constructions and the definitions.
Let $V=i_{*}\langle\widehat{\gamma}_{1},...,\widehat{\gamma}_{n}\rangle\leq H_{1}(\widehat{S};\mathbb{R})$ be as in Proposition 5.7 so that $\text{rank}(\mathcal{L})\geq\dim_{\mathbb{R}}V$ holds. Proposition 5.10 computes $\dim_{\mathbb{R}}V=\#\gamma-N_{co}$ when $(\gamma,\underline{\alpha})$ is not jointly orientable (i.e., when $\varepsilon=-1$), and Lemma 5.3 computes $\widehat{g}/2+1=g+\frac{1}{4}\#\underline{\kappa}_{\operatorname{odd}}+\frac{1}{2}$. Corollary 5.4 applies and we conclude that $\mathcal{L}=\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};\varepsilon)$. ∎
We now specialize this to the setting of Conjecture 1.1. Given a pants decomposition $\gamma$ of a topological surface $S$ and a length vector $\bm{\ell}\in\mathbb{R}_{>0}^{3g-3}$, let $\nabla_{\gamma}(\bm{\ell})$ denote the number of components of $S\setminus\gamma$ that have property ()($\nabla$)( ∇ ) from the introduction, i.e., such that the lengths of two of its boundaries sum to the length of the third (this exactly corresponds to the spine having a 4-valent vertex and $\underline{\alpha}$ being co-orientable on that component).
Corollary 5.12.
For any pants decomposition $\gamma$, any $\bm{\ell}\in\mathbb{R}_{>0}^{3g-3}$, and any $\mathbf{h}\in\mathbb{R}_{>0}^{3g-3}$ satisfying $\bm{\ell}\cdot\mathbf{h}=1$, set
$\mathcal{L}:=\overline{P\cdot\mathcal{Q}\mathbb{T}_{\gamma}(\bm{\ell},\mathbf{h})}.$
  1. (1)
    If $\nabla_{\gamma}(\bm{\ell})\leq 2g-5$, then $\mathcal{L}$ is the (connected) stratum $\mathcal{Q}^{1}\mathcal{M}_{g}(1^{4g-4-2\nabla_{\gamma}(\bm{\ell})},2^{\nabla_{\gamma}(\bm{\ell})};-1)$.
  2. (2)
    If $\nabla_{\gamma}(\bm{\ell})=2g-2$ and if $(\gamma,\underline{\alpha}(X))$ is jointly orientable for some (any) $X\in\mathbb{T}_{\gamma}(\bm{\ell})$, then $\mathcal{L}$ is a component of $\mathcal{Q}^{1}\mathcal{M}_{g}(2^{2g-2};+1)$.
@@
The first statement follows immediately from Corollary 5.11. For the second, we note that in this case the corresponding twist torus lives in a stratum consisting of squares of abelian differentials. Since each component of $S\setminus\gamma$ has genus 00, the curves of $\gamma$ span a $g$-dimensional subspace of $H_{1}(S)$, hence $\mathcal{L}$ has rank $g$. Work of Mirzakhani and Wright [MW17] classifies rank $g$ AISs in strata of abelian differentials as either the entire stratum component or the hyperelliptic locus. One then checks by inspection that if a twist torus lives entirely in the hyperelliptic locus then each of its curves must be fixed by the hyperelliptic involution, which is impossible when $\gamma$ is a pants decomposition. ∎
We now assemble the proof of (a more detailed version of) our main theorem.
Proof of Theorem 1.2.
Let $\gamma$ be any pants decomposition and fix any $\bm{\ell}$ and $\mathbf{h}$ as in the statement of the Theorem. Since pants are determined by their boundaries, $\bm{\ell}$ specifies a hyperbolic structure on $S\setminus\gamma$; let $\mathsf{Y}$ denote its spine and $\underline{\alpha}$ the dual arc system. We note that the edge lengths of $\mathsf{Y}$ are a linear function of the lengths $\bm{\ell}$ (see [CF24b, Figure 1]), so it is equivalent to rescale $\bm{\ell}$ and to rescale $\mathsf{Y}$.
Form the twist tori $\mathcal{P}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ and $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ as in §§3.1 and 4.1, as well as their associated measures $\mu_{\gamma}(\mathsf{Y},\mathbf{h})$ and $\nu_{\gamma}(\mathsf{Y},\mathbf{h})$. Recall that $\mathcal{O}$ interchanges these tori and that $g_{t}$ takes twist tori to twist tori (5):
The spine $\mathsf{Y}$ has $22g-$ 2 components, $\nabla_{\gamma}(\bm{\ell})$ of which have a unique 4-valent vertex and the rest of which have two trivalent vertices. By Theorem 2.3 (3), we can read off the stratum containing the flat twist torus:
$\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})\subset\mathcal{Q}^{1}\mathcal{M}_{g}(1^{4g-4-2\nabla_{\gamma}(\bm{\ell})},2^{\nabla_{\gamma}(\bm{\ell})};\varepsilon)$
where $\varepsilon=+1$ if $(\gamma,\underline{\alpha})$ is jointly orientable and $-1$ otherwise.
By Corollary 4.7 and (5), we have that these flat twist tori equidistribute outside a set $Z$ of zero density:
[…]
where $\mathcal{L}$ is the $P$-orbit closure of $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$. Applying Theorem 1.4, we therefore get that
[…]
Since $\nu_{\mathcal{L}}$ lives in the same ambient stratum as $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$, this proves part (2) of Theorem 1.2.
Of course, we can do even better and identify the orbit closure in most cases. Corollary 5.12 tells us that if $\nabla_{\gamma}(\bm{\ell})\leq 2g-5$, or $\nabla_{\gamma}(\bm{\ell})=2g-2$ and we have joint orientability, then $\mathcal{L}$ is the entire ambient stratum component. In particular, if $\nabla_{\gamma}(\bm{\ell})=0$ we have that $\mathcal{L}$ is the principal stratum, and so
[…]
where the last equality comes from Theorem 2.3 (2). This completes the proof of the Theorem. ∎
5.4. Other high rank twist tori
Here are some more concrete families of hyperbolic surfaces where the same argument applies. The proofs of the following results are identical to the proof of Theorem 1.2 (apart from some arithmetic to check that one can apply Corollary 5.11) and are therefore left to the reader.
In the following examples, the limiting distributions on $\mathcal{P}^{1}\mathcal{M}_{g}$ are qualitatively different from Mirzakhani measure. For example, the generic surface can contain large embedded disks, can be hyperelliptic, can have constrained thick-thin decomposition, and so on. We have chosen these families to be illustrative, not exhaustive — the reader is encouraged to come up with their own.
Regular plumbing fixtures
For $p\geq 3$, a regular plumbing fixture is a hyperbolic metric on a $p$-holed sphere with totally geodesic boundary and a dihedral symmetry group of order $2p$; see Figure 7. Regular plumbing fixtures with $p$ holes are parameterized by the length $L$ of each of their boundaries.
Figure 7. A regular plumbing fixture on a $7$-holed sphere.
Suppose that $\gamma$ is a multicurve on $S$ such that each component $\Sigma_{1},\ldots,\Sigma_{k}$ of $S\setminus\gamma$ has genus 0; let $p_{i}$ denote the number of boundary components of $\Sigma_{i}$. Taking each component to be a regular plumbing fixture with boundary lengths $L$, we get a twist torus whose corresponding metric ribbon graph $\mathsf{X}=\mathsf{X}(p_{1},\ldots,p_{k})$ has $k$ components, the $i^{\text{th}}$ of which has two vertices of valence $p_{i}$ and edges of length $L/2$.
Theorem 5.13.
Let all notation be as in the previous paragraphs and fix any $\mathbf{h}\in\mathbb{R}_{>0}^{\#\gamma}$ such that $\bm{\ell}\cdot\mathbf{h}=1$. Then if at least 3 of the $p_{i}$ are odd, the twist tori $\mathcal{P}\mathbb{T}_{\gamma}(e^{t}\mathsf{X},e^{-t}\mathbf{h})$ equidistribute in $\mathcal{P}^{1}\mathcal{M}_{g}$ outside of set $Z$ of zero density to the pullback of Masur–Smillie–Veech measure on the corresponding stratum. That is,
[…]
where $\nu_{\mathcal{Q}}$ is the Masur–Smillie–Veech measure on $\mathcal{Q}=\mathcal{Q}^{1}\mathcal{M}_{g}\left((p_{1}-2)^{2},\ldots,(p_{k}-2)^{2};-1\right)$.
We note that if at least $3$ of the $p_{i}$ are odd (which is what allows us to apply Corollary 5.4), then by the classification of strata of quadratic differentials [Lan08, CM14] we know that $\mathcal{Q}$ must be connected.
As in Corollary 5.12, we can also deduce a complementary result when the twist torus lands in a stratum of squares of abelian differentials. However, in this case the gluing pattern determines the limiting measure.
Theorem 5.14.
If all $p_{i}$ are even and $(\gamma,\underline{\alpha})$ is jointly orientable, where $\underline{\alpha}$ is dual to $\mathsf{X}$, then
[…]
If $S$ admits an involution fixing each of the curves of $\gamma$ setwise but reversing their orientations (hence $k\leq 2$), then $\mathcal{L}$ is the hyperelliptic locus in the component of the stratum $\mathcal{Q}^{1}\mathcal{M}_{g}\left((p_{1}-2)^{2},\ldots,(p_{k}-2)^{2};+1\right)$ that contains $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{X},\mathbf{h})$. Otherwise, $\mathcal{L}$ is the entire component.
Remark 5.15.
If $S$ admits a hyperelliptic involution $\iota$ that fixes $\gamma$ as a multicurve (but not necessarily each of its components), then one can consider the maximal $\iota$-invariant subtorus inside of $\mathbb{T}_{\gamma}(e^{t}\mathsf{X})$. The same reasoning as above (combined with Remark 4.4) shows that these invariant subtori also equidistribute to the measure supported on the hyperelliptic locus in the ambient stratum.
Minimal strata
We can also construct flat twist tori whose $P$-orbit closures are different components of the minimal stratum of (squares of) abelian differentials. For example, let $\gamma$ be one of the multicurves shown in Figure 8. Its complement is a $2g$-holed sphere, and setting the weight of each blue arc equal to 1 determines a hyperbolic structure on $S\setminus\gamma$ all of whose boundaries have length $2$ except for one which has length $1$. The dual ribbon graph $\mathsf{X}$ has a unique vertex of valence $4g$ and all edges of length $1$. Moreover, it is easy to check by hand that the pair $(\gamma,\underline{\alpha})$ is jointly orientable and one can even read off the parity of the induced spin structure from the data of $\gamma$ and $\underline{\alpha}$ (see [CS21, §6.3]).
Figure 8. Twist tori corresponding to the minimal stratum of squares of abelian differentials. The leftmost torus lands in the hyperelliptic locus, while the other two land in the odd and even spin components (which diagram corresponds to which spin depends on the parity of $g$ mod $4$).
Theorem 5.16.
With $\gamma$ and $\mathsf{X}$ as in Figure 8 and any choice of $\mathbf{h}\in\mathbb{R}_{>0}^{\#\gamma}$ such that $\bm{\ell}\cdot\mathbf{h}=1$,
[…]
where $\mathcal{Q}$ is the component of the stratum $\mathcal{Q}^{1}\mathcal{M}_{g}(4g-4;+1)$ specified by the gluing pattern of $\gamma$.

### v3

@@ 6.3. Lifting homologies
Let $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})\subset\mathcal{QM}_{g}(\underline{\kappa};-1)$ be a flat twist torus parametrizing quadratic differentials which are not squares of abelian differentials. If
$\gamma\geq g+\frac{1}{4}\#\underline{\kappa}_{\operatorname{odd}}+N_{o}(\mathsf{Y})+\frac{1}{2},$
then the only AIS containing $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is the ambient stratum component.
Applying this to our motivating example, we get that twist tori constructed out of regular plumbing fixtures generally equidistribute to the pullbacks of Masur–Smillie–Veech probability measures. We note at least $3$ of the $p_{i}$ must be odd in order to have high rank (see Remark 6.6), in which case we know that $\mathcal{Q}$ must be connected [Lan08, CM14].
In the following two results, the proof follows along the same lines as in §5.2, we just check that we can apply Corollary 6.8.
Theorem 6.9.
Consider any gluing of regular plumbing fixtures with boundary lengths $L$, as in Section 6.1. Suppose that at least 3 of the $p_{i}$ are odd and that the pair $(\gamma,\mathsf{X})$ is not jointly orientable. Fix any choice of heights $\mathbf{h}$ such that $\sum h_{i}=1/L$. There is a set of times $Z$ of zero density such that
[…]
where $\nu_{\mathcal{Q}}^{1}$ is the Masur–Smillie–Veech measure on $\mathcal{Q}^{1}\mathcal{M}_{g}\left((p_{1}-2)^{2},\ldots,(p_{k}-2)^{2};-1\right)$.
@@
Let $p_{\text{odd}}$ and $p_{\text{even}}=k-p_{\text{odd}}$ denote the number of $p_{i}$ which are odd and even, respectively. Then:
$\gamma=\frac{1}{2}\sum_{i=1}^{k}p_{i},\qquad\#\underline{\kappa}_{\operatorname{odd}}=2p_{\text{odd}},\qquad\text{and}\qquad N_{o}(\mathsf{Y})=p_{\text{even}}.$
We also compute that $4g-4=2\sum_{i=1}^{k}(p_{i}-2)$, and so we satisfy the hypotheses of the Corollary if and only if
$\gamma\geq(\#\gamma-k+1)+\frac{1}{2}p_{\text{odd}}+p_{\text{even}}+\frac{1}{2},$
which by arithmetic occurs if and only if $p_{\text{odd}}\geq 3$. ∎
We conclude with a substantial generalization of Theorem 1.2 which uses far fewer curves.
Theorem 6.10.
Suppose that $\gamma$ is a multicurve on a surface $S$ with $\gamma\geq 2g$. Fix a trivalent $\mathsf{Y}\in\mathcal{TRG}(S\setminus\gamma)$ and pick any $\mathbf{h}$ such that $\bm{\ell}\cdot\mathbf{h}=1$. There is a set of times $Z$ of zero density such that
[…]
In particular, projecting down to $\mathcal{M}_{g}$, the twist tori $\mathbb{T}_{\gamma}(e^{t}\mathsf{Y})$ for $t\not\in Z$ equidistribute to $\frac{B(X)}{b_{g}}d\operatorname{vol}_{\operatorname{WP}}(X)$.
@@
The condition that $\mathsf{Y}$ is trivalent gives that $\underline{\kappa}_{\text{odd}}=4g-4$, and since no trivalent ribbon graph is orientable, $N_{o}(\mathsf{Y})=0$. Therefore, we have
$g+\frac{1}{4}\#\underline{\kappa}_{\text{odd}}+N_{o}(\mathsf{Y})+\frac{1}{2}=2g-\frac{1}{2}.$
Thus, if $\gamma\geq 2g$, Corollary 6.8 identifies the $P$-orbit closure of the corresponding flat twist torus. ∎